# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 4.2.1 Introduction of Causal Representation Learning
To conquer the problem, causal representation learning tries to combine the advantages of the two sub-fields to extract the important features and the relationships with causality behind data automatically Schölkopf et al. (2021).
Figure 16: Illustration for the workflow of a typical causal representation learning agent. In the top panel, we have a depiction of a real block world with a robotic arm, along with the corresponding causal graph representing the relationships between the variables in the real world, nevertheless, it is totally unknown for the agent. Additionally, we have the data in the form of images generated by the block world. In the bottom panel, we see the structure of the causal representation learning agent. It utilizes a convolutional neural network (CNN) to extract meaningful representations from the data and perform downward tasks based on these representations. The encoder plays a crucial role in capturing relevant features and patterns from the input data, enabling the agent to learn and understand causal relationships.
@@
A typical scenario of causal representation learning can be depicted by Figure 16. Suppose there are a set of variables and causal mechanisms (including causal graphs and structural equations) between them which can describe how the world of robotic arms and colorful blocks works, whereas, they are not observable for the agent of causal representation learning. What the agent can observe is a set of images which are generated by robotic arms and blocks. The objective of a causal learning framework is to extract the causal variables and mechanisms from the observed images, and the variables and mechanisms can be used to implement other downstream tasks.
Formally, if the input data can be described by $X=(\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{p})$, causal representation learning will construct two functions and a directed acyclic graph, which include a causal representation function: $\tilde{\phi}:\mathcal{R}^{q}\rightarrow\mathcal{R}^{p}$, such that:
$X=\tilde{\phi}(Y),$ (49)
where, $Y=(\boldsymbol{y}_{1},\boldsymbol{y}_{2},...,\boldsymbol{y}_{q})$ is a set of causal variables and $q\ll p$; a directed acyclic causal graph $\mathcal{G}_{S}\equiv(Y,E)$, on which the node set is $Y$, and the directed links in $E$ represent causal relations; as well as a set of causal mechanisms:
$\boldsymbol{y}_{i}=f_{i}(PA_{i},U_{i}),$ (50)
for $i=1,2,…,q$, where $f_{i}$ is the structural equation for the $i$th causal variable $\boldsymbol{y}_{i}$, which describes how the parent causal variables $PA_{i}\equiv\{j|j\rightarrow i\in E\}\subset\{\boldsymbol{y}_{i}\}$ and the external unknown variable $U_{i}$ determine the value of $\boldsymbol{y}_{i}$. And there is a requirement for $U_{i}$ for all $i,j$, that is:
$P(U_{i},U_{j})=P(U_{i})P(U_{j}).$ (51)
Thus, any two external unknown variables need to be independent each other.
However, learning the three elements including the two functions and a graph in the same time is very hard, we can convert the problem to reduce the difficulty. This is to learn disentanglement representations, in which the representations are independent(disentangled) each other Wang et al. (2022). We will show how we can derive disentanglement representations from Equations 49 to 51.
Notice that if we combine the equations $\boldsymbol{y}_{i}=f_{i}(PA_{i},U_{i})$ for all $i\in[1,q]$ together, we have $Y=F(Y,U)$, where $U=(U_{1},U_{2},...,U_{n})$, and $F=(f_{1},f_{2},...,f_{q})$. That means, the causal mechanisms actually describe how the unknown variables $U$ determine the causal variables $Y$. Suppose we can solve this equation to obtain $Y$ expressed by $U$, and we define $\tilde{F}$ such that: $Y=\tilde{F}(U)$, according to Equation 49,
$X=(\tilde{\phi}\circ\tilde{F})(U)$ (52)
if we define $\phi\equiv\tilde{\phi}\circ\tilde{F}$, then $\phi$ can be understood as a decoder which can decode a disentanglement representation to data. Further, if we could construct a encoding function $\phi^{{\dagger}}:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ which takes $X$ to $U$, and we require that:
$\phi\circ\phi^{{\dagger}}\sim I,$ (53)
where, $I$ is the identity mapping, thus, $\phi\circ\phi^{{\dagger}}$ is an auto-encoder, which can encode the input data $X$ into a latent space $U$ and reconstruct $X$ according to $U$. While, according to Equation 51, the latent variables are independent each other. That is, we represent the input data by a set of independent or disentanglement latent variables, and we can reconstruct the input data according to these independent variables. Therefore, this is called disentanglement representation learning. In practice, independent component decomposition method can be use to construct the decoder Hyvarinen and Morioka (2016); Khemakhem et al. (2020). To reduce the computational complexity, we can also use invertible neural network to model $\phi$ Dinh et al. (2014).
With the disentangled representation variables $U_{i}$, we can effectively perform downstream tasks. One prominent application of disentangled representation learning is causal discovery. By leveraging the disentangled representations, we can uncover causal relationships between variables and gain insights into the underlying causal structure of a system Monti et al. (2020). Causal discovery allows us to identify cause-and-effect relationships and understand how different factors interact and influence each other. This knowledge can be invaluable in various domains, such as understanding complex systems, making informed decisions, and designing interventions or policies.
To conclude, a typical causal machine learning or disentanglement representation learning framework is always composed of a module for causal or disentanglement representations and a module which contains sub-units for different downstream tasks. The representation module, which contains an encoder and a decoder, tries to extract abstract representations of the data, and the representations are not for prediction but for causal inference. The downstream tasks may be variant. For example, the transferable mechanism learning tries to look for independent causal mechanisms to complete prediction tasks in very different environments Peters et al. (2016). And existing studies have shown that learning causal models may help in transferring modules across different domains Cui and Athey (2022); Kuang et al. (2018). Causal representation learning has been applied to various problems.
4.2.2 Comparisons between Causal Emergence Identification and Causal Representation Learning
There are similarities and differences between causal emergence and causal representation learning (CRL). We will introduce them in detail separately.
Similarities
@@ 4.2.2 Comparisons between Causal Emergence Identification and Causal Representation Learning
It is interesting to compare the tasks of causal emergence identification under the theoretic framework of Erik Hoel and causal representation learning. Actually, the task of causal emergence identification can be understood as a task of causal representation learning, where the macro variables are the causal variables, the macro dynamics are the causal mechanisms, the coarse-graining strategy is an encoding process from the original data to the representations. Under this framework, EI can be understood as a measure of the strength of the causal effect on the mechanism. To identify causal emergence, we need to learn the appropriate causal variables to represent data and to discover causal mechanism on the macro-level (on representational latent space) such that the EI of the learned causal mechanism in latent space is larger than the original data. We use Table 1 to clarify the comparisons item by item.
Table 1: Comparison between causal emergence identification and causal representation learning.
@@
Data Observations (time series) of micro-states Raw data generated by some causal mechanism in real
@@
Objective Finding an optimal coarse-graining strategy and a macro-dynamic that has stronger causal effect Finding an optimal representation of the raw data such that the independent causal mechanism can be realized by the representations.
@@
On the other hand, the learned abstract causal representations can be viewed as macro-states, which enhances the explainability of causal representation learning. This perspective allows us to consider that CRL is actually uncovering the hidden causal emergent features within the data.
@@
Moreover, these similarities between emergent phenomena and CRL contribute to a deeper understanding of emergence itself. By applying CRL frameworks to data generated by dynamical systems exhibiting emergent phenomena, we can extract a more profound causal structures. These profound causal structures may serve as the origins of the mysteries surrounding emergence. By delving into the depths of CRL and uncovering the hidden causal relationships within complex systems, we gain insights into the mechanisms behind emergent phenomena. These causal structures provide a foundation for understanding the emergence of novel properties and behaviors that arise from the interactions of simpler components. Exploring these intricate causal structures brings us closer to unraveling the enigma of emergence and sheds light on the fundamental principles governing complex systems. Consequently, the identification of emergent phenomena requires an observer, whether it be a human or a machine, to actively search for them.
By employing a machine observer, it is possible to potentially resolve the longstanding debates regarding whether emergent causality is epistemic or ontological. This is because the machine observer can provide a relatively objective outcome, although it is important to note that the result is still influenced by the observer. Furthermore, an intriguing area for future exploration involves investigating how the structure and capabilities of the observer can impact the emergence results that are identified by modifying the parameters of machine learning. This highlights the potential influence of the observer’s characteristics on the outcomes obtained through the rewiring of machine learning parameters.
Differences
@@
However, there are several differences between the CRL and causal emergence in theory. The biggest gap is that CRL assumes that there is a real causal mechanism behind it, and the data is generated by this causal mechanism Schölkopf et al. (2021). While for causal emergence identification, the emergent variables and mechanisms on macro-level are just handy ways to observe and understand data, "real causality" may not exist. The difference can be illustrated by Figure 17
@@
Figure 17: The Comparison between Causal Emergence Identification and Causal Representation Learning. All the potential co-founders are ignored. The black arrows represent causal relations, while the blue arrows represent interactions between macro-variables. The dashed directed lines represent the information flows for coarse-graining, representation or data generation.
@@
While, if epistemological perspective is adopted, this difference disappears because all they do is to extract meaningful information from observational data to obtain representations with stronger causal effects.
@@
Another major difference is that the causal mechanism for macro-states in causal emergence is a dynamical system. Therefore, if the dynamics in macro-state space is on a network, circular structures may exist because the state variables will be iterated in the dynamical system. As a result, feed-backs and circular indirect interactions are allowed in such model. However, the causal mechanism is always represented by a structural causal model in CRL. While, these models are always directed acyclic graphs, therefore, loops or circular structures are not allowed. The problem is that the meaning of the links in a network dynamics model is different from the links in structural causal model. Hence, the conflict can be resolved if we comprehend that the links in a structural causal model depict an average (net) causal effect between variables when they are in a steady state. On the other hand, the links in network dynamical models within causal emergence represent instantaneous interactions.
@@
Finally, unknown co-founders are always ignored in causal emergence, while, they play very important roles in causal structural models.
4.2.3 Out-of-distribution Generalization
Causal representation learning is established in a self-supervised manner, in the sense that it reveals the intrinsic causal structure of data. Thus, the learned causal representations could not be exploited directly for supervised prediction task, where the causation between data and a specified target must be characterized. The measurement of causal emergence by vanilla Effective Information Hoel et al. (2013) is also defined for stationary dynamics systems in a self-supervised fashion, since the input(cause) space and the output(effect) space share the support. However, Hoel generalizes Effective Information to general input-output systems Chvykov and Hoel (2020)(see section 3.6.1 where the cause and the effect space take arbitrary support, which could either be discrete or continuous.
Empowered by the general notation of EI, causal emergence could be applied in supervised learning to evaluate the strength of causation between the feature space $\mathcal{X}$ (or a representation learned upon which) and the target space $\mathcal{Y}$, thus enhancing the prediction from the cause(feature) to the effect(target). It is worth noting that a direct fitting from $\mathcal{X}$ to $\mathcal{Y}$ on the observations suffice for common prediction tasks with an i.i.d. assumption, implying that the training and test data are independently and identically distributed.
However, if samples are drawn out of the training distribution, it is essential to learn a representation space that generalizes from training to test environment, known as out-of-distribution (OOD) generalization. Due to the common belief that causal relations generalize better than statistical correlation Arjovsky et al. (2019), causal emergence theory could serve as a criteria for the causation embedded in the representation space. The occurrence of causal emergence indicates the revelation of potential causal factors of the target, thus producing a robust representation space with regarding to out-of-distribution generalization.
Introduction of OOD
Before the detailed discussion of the connection between causal emergence and OOD generalization, we begin with a brief introduction to the latter topic. Out-of-distribution generalization towards arbitrary distributional shifts is generally impossible Ye et al. (2021). Therefore, it is commonly assumed there exists an underlying causal mechanism that governs the data generation process, which remains invariant to distributional shifts. Formally, by Pearl’s Structural Equation Model Pearl (2009) which causally formulates data generation, the target $Y$ is supposed to be generated from part of the feature $X$, i.e. $Y=f(X_{c},\epsilon)$ where $X_{c}$ is selected features or a representation of $X$, and $\epsilon$ is a random noise independent of $(X,Y)$. Obviously $X_{c}$ is the true cause of the target, and the causal mechanism $f$ is presumed to be invariant across training and test environments. On the other hand, the raw feature $X$ might be a mixture of causal variables and non-causal variables. The latter could be produced by confounding, selection bias or anti-causal mechanisms, all of which introduce spurious correlation between the target and non-causal variables. The strength and direction of spurious correlation are variant across environments, resulting in distributional shifts that hampers out-of-distribution performance of machine learning techniques. As a remedy, a prominent branch of OOD algorithms Arjovsky et al. (2019); Huh and Baidya (2022); Zhao et al. (2019); Ganin et al. (2016) learn an invariant representation space distilling the causal variables from the raw feature space. Typically a metric of the representation is devised to verify the stability of correlation between the learned representation space and the target, and is adopted as a regularization for the optimization objective to encourage the representation learner to extract causal variables from features. The paradigm for invariant representation learning could be formulated as Wang et al. (2022):
where $\phi(X)$ is the learned representation, $f$ is the learned predictor, $\mathcal{R}$ is the task-specific risk function and $\ell_{reg}$ is the regularization.
Representation learning based OOD algorithms are diversified in their constraint of the representation space characterized by the regularizers. Each metric of the representation fits a specific data generation pattern, and does not guarantee to filter out the causal variables when the setting is violated. However, causal emergence potentially offers a united metric of representation for OOD generalization rooted from causality theory.
Before diving into the application of causal emergence, we would first discuss representative regularizers of representation learning and their limitations. Domain invariant representation learning proposes an aligned feature space with the constraint $g(X)\perp\mathcal{E}$, where $\mathcal{E}$ denotes the environment index for training and test environments. The regularization is implemented by the metric of Maximum Mean Discrepancy (MMD) distance of representation across environments Li et al. (2018) or an adversarial classifier for the environment index Ganin et al. (2016). However, the alignment constraint provides generalization guarantee only for covariate shift Zhao et al. (2019) where the distribution shift is restricted to the features $X$ without perturbing $P(Y|X)$. On the other hand, invariant risk minimization Arjovsky et al. (2019); Koyama and Yamaguchi (2020) pursues an invariant conditional distribution $P(Y|g(X))$ of the representation across environments, which translates to $Y\perp\mathcal{E}\;|\;g(X)$. Invariant risk minimization is proven to recover the causal variable in a data generation process where the non-causal variable is anti-causally generated by the target, but is reported to fail under covariate shift Ahuja et al. (2021). Further, conditional invariant representation Li et al. (2018); Huh and Baidya (2022) is designed specifically for image classification where the target is modeled as the cause of the feature instead of the effect. The conditional independence $g(X)\perp\mathcal{E}\;|\;Y$ is enforced on the representation space.
Actually, it is theoretically proven that any statistical independence regularization is valid for a subgroup of Structural Equation Models and it does not generalize across all data generation processes Kaur et al. (2022). As a result, the metric of representation manages to capture the causation between the representation space and the target variable only if a specific causal graph of data generation is given, which hinders the general applicability of invariant representation learning. Worse still, the existence of the environment index $\mathcal{E}$ in the independence regularizers introduces the necessity of multiple environments for training procedure, increasing the data collection and labelling expenses. To this end, there is active demand for a unified and general metric of representation for OOD generalization, for which causal emergence might be a good candidate.
Causal emergence and OOD
The match between OOD generalization and causal emergence is attributed to the do intervention of Effective Information. By definition Pearl (2009), intervening the feature distribution with $do(X)$ modifies the structural equations of all the variables in feature space to a constant value while the equation for the target remains intact as $Y=f(X_{c})$. On the one hand, the assumption of causal mechanism’s invariance to distributional shift is respected, since the causal correlation between causal variables and the target is reserved. On the other hand, the spurious correlation between non-causal variables and the target are removed by the intervention. For example, a non-causal variable anti-causally correlated with the target leads to a structural equation as $X_{n}=h(Y,\epsilon)$, and $do(X_{n})$ replaces the equation by $X_{n}=U$, where $U\sim U_{max}$, and $U_{max}$ represents the maximum entropy distribution on the domain of $X_{n}$. As a result, $X_{n}$ is no longer caused by $Y$ and the spurious correlation vanishes. For another example, if a non-causal variable is correlated with the target by a confounder $E$ (probably in the feature space), the structural equation takes the form of $X_{n}=\phi(E,\epsilon)$ and $Y=f(X_{c},E,\epsilon)$. Then the do intervention would modify the equation to $X_{n}=U$ and confounding ceases to exist. In summary, the intervention introduced by EI captures the causal correlation between features and the target whereas suppresses the spurious correlation. It renders EI an ideal metric to assess the causality contained by the representation space and thus the generalization of the induced algorithm.
Recall that an increment in the Effective Information implies the occurrence of causal emergence. Therefore, we bring forward a conjecture that out-of-distribution generalization could be achieved simultaneously with maximizing EI (or normalized EI) for representation learning. Notably, Shannon’s mutual information holds the property that any representation has no gain of mutual information with the target over the raw feature space, i.e. $I(g(X);Y)\leq I(X;Y)$, by data processing theorem. Thus, direct fitting inclines to absorb as much information with the target as possible into the representation space. In contrast, EI(or normalized EI) would possibly peak at a medium stage of abstraction of raw feature, coinciding with the philosophy of OOD generalization that less could be more. Ideally, at the peak of EI(or normalized EI) where causal emergence occurs, all non-causal features are excluded and the causal features are revealed, resulting in the most informative representation while performing invariant to distributional shifts.
Actually, removing spurious correlation and recovering causal mechanisms by manipulating feature distributions is shared by the reweighting technique widely adopted in OOD literature. For instance, Stable Learning Shen et al. (2020a, b); Zhang et al. (2021) is oriented at the setting where collinearity exists in input variables, causing spurious correlation between non-causal variables and target. It reweights training samples to decorrelate the features and thus reducing collinearity and eliminating spurious correlation. Similar to EI, Stable Learning is free of environment index labeling. Further, sample reweighting and feature decorrelation shares the philosophy of distribution intervention. In this sense, EI could be viewed as an information-theoretic abstraction of reweighting-based debiasing techniques for OOD generalization.
4.3 Other Related Topics in AI
There are other related topics in artificial intelligence, e.g., abstraction of causal models and world model based reinforcement learning. We will introduce them in detail.
4.3.1 Abstraction of Causal Models
In CRL, we have the ability to extract causal relationships and construct causal models from raw data. However, in cases where the causal model becomes excessively large and complex, it becomes necessary to employ an abstraction process Rubenstein et al. (2017). Abstraction in causal modeling refers to the reduction or simplification of the causal model while preserving its essential causal relationships and important features.
Coarse-graining is the multi-scale description of the state information on different information reduction scales. While causal models also can be multi-scale Chalupka et al. (2016). The causal model in the higher level is an abstraction of the model in the lower level.
If the original causal model is given, then the key problem is how can we obtain an abstraction of the original model. Obviously, the abstraction of the causal model cannot be arbitrary, it must recover the main features of the original causal model. Several works proposed different standards for abstract models Rubenstein et al. (2017); Beckers and Halpern (2019); Chalupka et al. (2016), among which approximate abstraction is a good and reasonable proposal. The basic idea behind the method of approximation abstraction is that the differences of causal mechanism both in the abstraction and the original causal model should be minimized Rubenstein et al. (2017); Beckers and Halpern (2019); Beckers et al. (2020); Beckers (2021).
Figure 18: An approximate abstraction of the El Niño phenomenon.
For example, we can build a causal model to describe how the Equatorial Pacific Zonal wind ${W}$ can have a causal effect on the temperature of the Equatorial Pacific ${T}$ Chalupka et al. (2016). We write the causal variable $W$ as $X_{C}$, the effect variable $T$ as $X_{E}$, and this is the original causal model, which can be denoted as $\mathcal{M}_{m}=((\{{U}\},\{{X_{C}},{X_{E}}\},\mathcal{R}_{m}),\{g_{m},f_{m}\},\mathcal{I}_{m})$, where ${U}$ is the collection of all the unknown variables, ${X_{C}}$ and ${X_{E}}$ are causal variables, $g_{m}$ and $f_{m}$ are causal mechanisms to describe how ${U}$ and ${X_{C}}$ can affect other variables, $\mathcal{I}_{m}$ is the set of interventions. However, we care about higher level causality when we consider the El Nino or La Nina phenomena. Thus, we can build an abstract causal model $\mathcal{M}_{M}=((\{{U}\},\{Y_{C},Y_{E}\},\mathcal{R}_{M}),\{g_{M},f_{M}\},\mathcal{I}_{M})$, where $Y_{C},Y_{E}$ are new causal variables which are built by the coarse-graining mapping $\phi$, that is $Y_{C}=\phi({X_{C}})$, $Y_{E}=\phi({X_{E}})$. And $g_{M}$ and $f_{M}$ are causal mechanisms of $Y_{C},Y_{E}$ on the abstraction model. However, if $\phi$ is given, $g_{M}$ and $f_{M}$ are limited otherwise inconsistencies will exist. In  Beckers et al. (2020), the authors propose a limitation on $g_{M}$ and $f_{M}$. We denote $d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})$ as the distance between different level models $\mathcal{M}_{m}$ and $\mathcal{M}_{M}$. They claim that $\mathcal{M}_{M}$ is an $\epsilon$-approximate abstraction of $\mathcal{M}_{m}$, if and only if $d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})$ must be less than $\epsilon$ Beckers et al. (2020). That is,
$d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})=\min_{\phi_{\mathcal{U}}}\max_{do(X=x)\in\mathcal{I}_{m},{u}_{m}\in\mathcal{R}_{m}(\mathcal{U}_{m})}$
$\big{\{}d_{v_{M}}(\phi(\mathcal{M}_{m}({u}_{m},do(X=x))),\mathcal{M}_{M}(\phi_{\mathcal{U}}({u}_{m}),\omega_{\phi}(do(X=x))))\big{\}}.$ (55)
In this equation, $\phi_{\mathcal{U}}$ is a coarse-grained mapping of unknown exogenous variables, $do(X=x)$ is an intervention on certain endogenous variables, $d_{v_{M}}$ is the distance between variables defined at a higher level, and $\omega_{\phi}$ is the mapping of interventions between different causal models at different levels. The distance between models is transformed into the distance between variables. Given a defined method of measuring the distance between variables (such as Euclidean distance), we can measure the distance between the output of a lower level model after coarse-grained projection$\big{(}\phi(\mathcal{M}_{m}({u}_{m},do(X=x)))\big{)}$ and the variables output by a higher level model$\big{(}\mathcal{M}_{M}(\phi_{\mathcal{U}}({u}_{m}),\omega_{\phi}(do(X=x)))\big{)}$. We focus on identifying external variables and intervention methods that maximize the distance, and we do not concern ourselves with the coarse-grained mapping between external variables. Therefore, we select the optimal $\phi_{\mathcal{U}}$ and require that it be surjective Beckers et al. (2020). This yields a distance value that characterizes the inconsistency between different levels of causal models.
There are similarities between causal emergence identification problem and abstraction of causal models. If the original causal mechanism can be understood as a micro-dynamics, then the abstracted mechanism can be regarded as a macro-dynamics. Therefore, finding the optimal $\phi$ and causal mechanisms $g_{M}$ and $f_{M}$ is equivalent to finding an optimal coarse-graining strategy and macro-dynamics. In the framework of NIS, we set limitations on coarse-graining strategy and macro-dynamics by requiring that the macro-dynamics can recover the micro-dynamics with $\epsilon$ error to exclude trivial cases. This requirement is also similar to the approximation requirement for the distance.
In artificial intelligence, model abstraction is a widely used technique to obtain a compressed model. For example, the abstraction of a neural network model can be stored in cell-phone to conduct fast computation Wang et al. (2021); Pulina and Tacchella (2010). Another article discusses the importance of structural analysis methods for neural network analysis and proposes a new method based on a formal theory of causal abstraction. This method aligns neural representations with variables in interpretable causal models and uses interchange interventions to experimentally verify their causal properties. The method is utilized in a case study to examine neural models that have been trained on a complex Natural Language Inference (NLI) dataset Geiger et al. (2021). This dataset comprises numerous artificially processed, linguistically intricate English phrases and sentences. The analysis reveals that neural representations effectively capture the compositional structure of the provided examples Geiger et al. (2021). Model reduction is also important topics in control theory and dynamical systems Antoulas (2005).
The primary distinction between causal emergence identification and causal model abstraction lies in their approach to abstraction or dimension reduction. Causal emergence identification focuses on abstracting or reducing dimensions from the original data, whereas causal model abstraction focuses on abstracting from the micro-level causal models. In another word, coarse-graining operates on states or data, while abstractions operate on models. In physics literatures Christensen and Moloney (2005), abstractions of models are also called renormalization.
Another difference is unknown confounder $\mathcal{U}$ being considered in causal model abstraction is ignored in all the discussion of causal emergence. The emergence of causality requires the construction of models from a multi-scale perspective, and how to discuss mixed variables at different scales has always been a challenge. The theory of approximate abstraction takes into account the coarse-grained mapping of latent confounding variables $\phi_{\mathcal{U}}$, making it compatible with the coarse-grained function of endogenous variables. For the study of causal emergence, this theory lays a theoretical foundation for considering the confounding effects brought by environmental factors.
The approximate abstract framework currently has many limitations. Firstly, there is a lack of effective algorithms that can calculate the degree of approximation in the approximate abstract. Given the involvement of traversing intervention sets, it is difficult for machines to perform efficient numerical calculations that are not maintainable, and cannot be used as constraint conditions for automatic modeling. Additionally, the intervention space is very large, and some interventions are meaningless or impossible, so it is necessary to empirically provide an intervention distribution in advance Beckers et al. (2020), making it difficult for this framework to be applied to real-world problems.
4.3.2 World Models in Reinforcement Learning
In both the generalized causal emergence identification problem and causal representation learning, a common computational framework is shared, consisting of an encoder-decoder architecture and causal mechanisms or dynamics. This framework can also be integrated into a reinforcement learning agent, enabling the agent to learn a world model and make informed decisions and plans. Such an agent, known as a world model-based reinforcement learning agent, has demonstrated impressive performance across various environments and tasks Ha and Schmidhuber (2018).
Reinforcement learning is a significant sub-field of machine learning that possesses several distinctive features when compared to conventional supervised learning techniques Sutton and Barto (2018). Firstly, reinforcement learning agents have the capability to interact with the environment and actively acquire new information through interventions. Consequently, causal information can be obtained through these interactions Zeng et al. (2023). Secondly, in contrast to supervised learning where the processes of learning and action are clearly separated, reinforcement learning agents can seamlessly blend learning and action processes. Lastly, reinforcement learning agents demonstrate strong performance in environments that have delayed feedback, such as game playing scenarios Mnih et al. (2013); Silver et al. (2016).
The reinforcement learning algorithms can be classified into model-based and model-free types, and the world model based reinforcement learning belongs to the former Ha and Schmidhuber (2018); Hafner et al. (2023); LeCun (2022); Hafner et al. (2019a, b, 2020, 2023). As its name suggests, there is a model of the world which can simulate the dynamics of the environment the agent confronts. The dynamics of the world model can be learned via the interactions between the agent and the environment, and it can also help the agent to make plans and decisions about the uncertain environment Sekar et al. (2020). However, in order to represent a complex environment, the world model needs to provide a coarse-grained description of the environment. Consequently, a standard architecture of a world model typically includes an encoder and a decoder, as illustrated in Figure 19.
Figure 19: The architecture of a world model embedded in a model-based reinforcement learning agent, where $X_{t}$ is the environment at time $t$, and $Y_{t}$ is its representation, $H_{t}$ is the internal variable of the agent. Different from NIS and causal representation learning agent, another variable, the action in the previous step $A_{t-1}$, should be input to the world model to implement the one step dynamics. And the world model can be trained in a supervised learning manner by comparing the predicted representation of the environment and the prediction.
The world model can also be treated as a macro-dynamics or causal mechanism Zhang et al. (2020); Allshire et al. (2021); Jiang et al. (2022); Zhang et al. (2020); Wang et al. (2022). All the states of the environment can be regarded as macro-states or causal representations. These states are constructed to ignore irrelevant information and capture the most important causal features of the environment for the agent to make decisions. In the process of planning, the agent can use the world model to simulate the running of the real world dynamics. Numeric results show that the reinforcement learning algorithms with world models can perform better than the classic reinforcement learning algorithms.
World model-based reinforcement learning techniques can also be employed for identifying causal emergence. When attempting to identify the causal emergence phenomenon within a highly complex system, extracting macro states and dynamics solely from observational data becomes challenging. Explorations or experiments are required to facilitate this process Sekar et al. (2020). In such cases, a reinforcement learning agent equipped with a trainable world model, specifically a macro-dynamics model, becomes essential. By treating the complex system as a unified entity, the agent with the world model can interact with it and derive emergent causal laws from these interactions.
On the other hand, the exploration of causal emergence can offer valuable insights into reinforcement learning algorithms. Firstly, the measure of causal effect, EI, can be applied to world models. Techniques of maximizing EI can be employed to enhance the causal effect of world models within reinforcement learning agents Yang et al. (2023). This, in turn, can greatly benefit planning and problem-solving tasks, improving their efficacy and performance. It is like the bi-simulation Kemertas and Aumentado-Armstrong (2021); Zhang et al. (2020) which can train the agent with an internal metric.
Second, if the reinforcement learning agent with a world model can learn the emergent causal mechanisms automatically from a complex system that the agent is interacting with, then it will have more complex behaviors and be more intelligent. The agent may use the learnt emergent causal mechanisms to make plans. From learning to planning, the whole process may take place continuously. Therefore, causal emergence can be observed from outside because new behaviors governed by different causal mechanisms emerge as time goes by. This agent seems to have its own purpose or objectives.
4.4 Summary

### 25 Feb 2024 v3

@@ 5.1 Causal Emergence and Causal Representation Learning
To address the problem, CRL tries to combine the advantages of the two sub-fields: representation learning and causal inference to extract the important features and the relationships with causation behind data automatically Schölkopf et al. (2021).
@@
A typical CRL scenario is depicted in Figure 16. Suppose there is a set of variables and causal mechanisms (including causal graphs and structural equations) that describes how the world of robotic arms and colorful blocks works. However, these mechanisms are not directly observable for the agent undergoing causal representation learning. Instead, what the agent can observe is a set of images generated by robotic arms and blocks. The objective of a causal learning framework is to extract the causal variables and mechanisms from the observed images, and the variables and mechanisms can be used to implement other downstream tasks, for example, causal relationship discovery Monti et al. (2020) and predictions in different environments Peters et al. (2016) or domains Cui and Athey (2022); Kuang et al. (2018).
Figure 16: Illustration of the workflow of a typical causal representation learning (CRL) agent. The top panel depicts a real block world with a robotic arm and the corresponding causal graph, representing relationships between variables in the real world, which is totally unknown to the agent. Additionally, the data are presented as images generated by the block world. The bottom panel depicts the structure of the causal representation learning agent. It utilizes a convolutional neural network (CNN) to extract meaningful representations from the data and perform downward tasks based on these representations. The encoder plays a crucial role in capturing relevant features and patterns from the input data, enabling the agent to learn and understand causal relationships.
When comparing Figures 14 and 15 with Figure 16, we found that as a framework for machine learning, Neural Information Squeezer Plus can be regarded as a kind of CRL framework. Thus, with the NIS and NIS+, we effectively employ CRL to address the problem of identifying causal emergence (CE). Both CE and CRL frameworks feature an encoder and decoder designed to represent raw data as causal variables. Latent causal mechanisms can be learned in both frameworks. In this section, we compare the similarities and differences between these two emerging fields.
5.1.1 Similarities
@@
It is interesting to compare the tasks of causal emergence identification within the theoretical framework of Erik Hoel and causal representation learning. Actually, the task of causal emergence identification can be understood as a task of causal representation learning, where the macro-variables are the causal variables, the macro-dynamics are the causal mechanisms, and the coarse-graining strategy is an encoding process that transforms the original data into representations. Within this framework, EI can be understood as a measure of the strength of the causal effect on the mechanism. To identify causal emergence, we need to learn the appropriate causal variables to represent data and discover causal mechanisms at the macro-level (in the representational latent space) to ensure that the EI of the learned causal mechanism in the latent space is larger than that of the original data. Table 1 shows a detailed comparison of causal emergence identification and causal representation learning.
Table 1: Comparison of causal emergence identification and causal representation learning.
@@
Data Observations (time series) of micro-states Raw data generated by some causal mechanism in real life
@@
Objective Finding an optimal coarse-graining strategy and a macro-dynamic that has a stronger causal effect Finding an optimal representation of the raw data to ensure that the independent causal mechanism can be realized by the representations
@@
On the other hand, the learned abstract causal representations can be viewed as macro-states, which enhances the explainability of causal representation learning. This perspective suggests that CRL essentially uncovers hidden causal emergent features within the data.
@@
Moreover, these similarities between emergent phenomena and CRL contribute to a deeper understanding of emergence itself. By applying CRL frameworks to data generated by dynamical systems exhibiting emergent phenomena, we can extract more profound causal structures. These profound causal structures may serve as the origins of the mysteries surrounding emergence. By delving into the depths of CRL and uncovering the hidden causal relationships within complex systems, we can gain insights into the mechanisms behind emergent phenomena. These causal structures may provide a foundation for understanding the emergence of novel properties and behaviors that arise from the interactions of simpler components.
5.1.2 Differences
@@
However, there are several theoretical differences between CRL and causal emergence. The biggest difference is that CRL assumes that there is a real causal mechanism behind the data, and the data are generated by this causal mechanism Schölkopf et al. (2021). In contrast, for causal emergence identification, the emergent variables and mechanisms at the macro-level are just handy ways to observe and understand data, and “real causality” may not exist. The differences are illustrated in Figure 17.
@@
Figure 17: Comparison of causal emergence identification and causal representation learning. All the potential co-founders are ignored. The black arrows represent causal relations, whereas the blue arrows represent interactions between macro-variables. The dashed lines with arrows represent the information flows for coarse-graining, representation, or data generation.
@@
However, if an epistemological perspective is adopted, this difference disappears because both approaches extract meaningful information from observational data to obtain representations with stronger causal effects.
@@
Another major difference is that the causal mechanism for macro-states in causal emergence is a dynamical system. Therefore, if the dynamics in the macro-state space are on a network, circular structures may exist because the state variables will be iterated in the dynamical system. As a result, feedback and circular indirect interactions are allowed in such a model. However, the causal mechanism is always represented by a structural causal model in CRL. Although these models are always directed acyclic graphs, loops or circular structures are not allowed. However, these differences are not important because Markovian dynamics can always be converted to causal models, as mentioned in Section 2.1.
@@
Finally, unknown co-founders are always ignored in causal emergence, whereas they play very important roles in causal structural models.